What are the first and second stage procedures in assigning costs to products when using an activity-based costing system? please provide a brief description of each.
Answer #1

First stage: In this stage of cost allocation the process is to determine the cost of each occurrence of an overhead event during the process. In this stage the expenses of certain overhead activities are divided to a per-event cost. For example suppose that a company manufactures wire ropes and has to reset the machines over two kinds of ropes that it makes. The cost of resetting was $100,000 and the number of resets was 1,000. Thus cost of one setting switch = $100,000/1000 = $100. Thus we can see that this stage is about allocation to activities.

Second stage: In this stage the process is to allocate the cost of each occurrence to the individual items that are being produced by the company. This stage entails allocating the activity cost to each product. Suppose that in the above example the company makes 100 units of rope type A or 200 units of rope type B after every resetting. So machine resetting/switching cost that will be allocated to a single unit of A will be = $100/100 units = $1 per unit. For B it will be = $100/200 units = $0.50 per unit. Thus we can see that this stage is about allocation to production.